# A torsion theoretic interpretation for sheaves of modules and Grothendieck topologies on directed categories

### v2

@@ Abstract.
We prove that every Grothendieck topology induces a hereditary torsion pair in the category of presheaves of modules on a ringed site, and obtain a homological characterization of sheaves of modules: a presheaf of modules is a sheaf of modules if and only if it is saturated with respect to torsion presheaves, or equivalently, it is right perpendicular to torsion presheaves in the sense of Geigle and Lenzing. We also study Grothendieck topologies on directed categories ${\mathscr{C}}$, and show that every Grothendieck topology on ${\mathscr{C}}$ is a subcategory topology if and only if ${\mathscr{C}}$ is an artinian EI category. Consequently, in this case every sheaf category is equivalent to the presheaf category over a full subcategory of ${\mathscr{C}}$. Finally, we classify all Grothendieck topologies on a special type of noetherian EI categories, and extend the locally self-injective property of representations of ${\mathrm{FI}}$ and ${\mathrm{VI}}$ to representations of their infinite full subcategories. Some potential applications in group representation theory are given at the end of this paper.
@@
Grothendieck topologies, sheaves, torsion pairs, sheafification, sheaf cohomology, Serre quotient, directed categories, EI categories
Z.X. Di was partly supported by NSF of China (Grant No. 11971388), the Scientific Research Funds of Huaqiao University (Grant No. 605-50Y22050) and the Fujian Alliance of Mathematics (Grant No. 2024SXLMMS04); L.P. Li was partly supported by NSF of China (Grant No. 12171146); L. Liang was partly supported by NSF of China (Grant No. 12271230).
@@ 1. Introduction
Sheaves of modules on ringed sites have been widely applied in algebraic geometry, algebraic topology, and geometric representation theory. Recently, they were also used to investigate modular representation theory of finite groups [3, 25, 26, 27] and continuous representations of topological groups [7]. This new approach combines ideas and methods from topos theory, representation theory of categories, and representation theory of groups, and hence not only proposes a comprehensive technique to study representations of groups, but also reveals deep relations among distinct areas. A classical example is Artin’s theorem [18, III. 9, Theorems 1 and 2], from which it follows that the category of continuous representations of a topological group $G$ is equivalent to the category of sheaves of modules over a ringed site whose underlying category is an orbit category of $G$ equipped with the atomic Grothendieck topology.
@@
The following result gives a characterization of sheaves of modules over $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ in terms of the torsion theory mentioned above, and extends [7, Theorem 3.8] to full generality. Compared to the original definition of sheaves in terms of amalgamations of matching families, this characterization is homological, and is more elementary for applications in representation theory.
@@
It is direct to check that $J$-saturated $\mathcal{O}$-modules are precisely $\mathcal{O}$-modules right perpendicular to $J$-torsion $\mathcal{O}$-modules in the sense of Geigle and Lenzing [12]; for details, see [12, Section 1]. Consequently, applying [12, Propositions 2.2 and 2.5], one immediately deduces the following result:
@@
It has been pointed out in [2, IV.9] by Grothendieck and Verdier that every Grothendick topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid (or called a subcategory topology) when ${\mathscr{C}}$ is a finite category, and in this case ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is equivalent to the category of presheaves of modules on a full subcategory of ${\mathscr{C}^{\mathrm{op}}}$. Similar results have been established by Hemelaer in [14] and Lindenhovius in [17] for posets satisfying certain extra conditions. Although it seems to be a very difficult task to classify small categories whose Grothendieck topologies are rigid, in this paper we obtain a complete answer for categories satisfying the following combinatorial condition: the binary relation $\leqslant$ on $\operatorname{Ob}({\mathscr{C}})$ defined by setting $x\leqslant y$ if and only if ${\mathscr{C}}(x,y)\neq\emptyset$ is a partial order. We call them directed categories. The reason we choose them is plain: many important categories of particular interest in representation theory are directed categories, for examples, posets, orbit categories of groups, quivers without loops or oriented cycles, and etc.
@@
Let ${\mathscr{C}}$ be a directed category. Then every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if ${\mathscr{C}}$ is a noetherian EI category; that is, the poset $(\operatorname{Ob}({\mathscr{C}}),\leqslant)$ satisfies the ascending chain condition and every endomorphism in ${\mathscr{C}}$ is an isomorphism.
Consequently, when ${\mathscr{C}}$ is a noetherian EI category, questions concerning ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ can be completely transferred to questions about representations of full subcategories of ${\mathscr{C}}$, so we have no “generic” sheaf theory in this situation. Explicitly, by the Comparison Lemma [13, C2.2, Theorem 2.2.3] (see also [13, C2.2, Example 2.2.4(d)]), one has:
@@
However, when ${\mathscr{C}}$ is not noetherian, a classification of Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ becomes much harder. In this paper we solve this question only for EI categories ${\mathscr{C}}$ of type $\mathbb{N}$ (see the beginning of Section 6 for a precise definition) or type $\mathbb{Z}$. Although it seems to the reader that too many combinatorial restrictions are imposed on them, there are still quite a few interesting examples whose representation theory are actively studied (see for instances [6, 20, 21]):
  • skeletal full subcategories of the category ${\mathrm{FI}}$ of finite sets and injections, where ${\mathrm{FI}}$ is equivalent to the opposite category of an orbit category of the permutation group on ${\mathbb{N}}$ ([13, Examples A2.1.11(h) and D3.4.10]);
  • skeletal full subcategories of the opposite category of the category $\mathrm{FS}$ of finite sets and surjections, where $\mathrm{FS}$ is equivalent to an orbit category of the automorphism group of the free Boolean algebra on a countable infinity of generators, or the self-homeomorphism group of the Cantor set ([13, Example D3.4.12]);
@@
We can use Theorem 1.6 and the Comparison Lemma to deduce some interesting results in representation theory of combinatorial categories. For instance, let $k$ be a field of characteristic 0, and let ${\mathscr{C}}$ be the category of finite sets and injections, or the category of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved in [11] (see also [22]) that every indecomposable projective representation of ${\mathscr{C}}$ over $k$ is also injective. It turns out this locally self-injective property holds for all (essentially) infinite full subcategories of ${\mathscr{C}}$.
@@
Let ${\mathscr{C}}$ be specified as above, $k$ a field of characteristic 0, and ${\mathscr{D}}$ a full subcategory of ${\mathscr{C}}$ containing infinitely many objects (up to isomorphism). Then every indecomposable projective ${\mathscr{D}}$-module over $k$ is also injective.
For a generalized version of this result, see Theorem 7.1.
@@ 2.1. Grothendieck topologies
Throughout this paper we fix $J$ to be a rule assigning to each object $x$ a collection $J(x)$ of left ideals of morphisms on $x$. The following definition is taken from [18, III.2, Definition 1], which defines a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ if $J$ satisfies all three axioms. Here we use the dual version, since we mainly consider representations of ${\mathscr{C}}$ which are covariant functors, rather than presheaves which are contravariant functors.
@@
The following fact is well known; see [18, Section III.2]. For the convenience of the reader, we give a complete proof.
@@ 2.2. Sheaf theory
Fix a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ and let $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ be the corresponded site. The following definitions are taken from [18, III.4].
@@
When $J$ is the trivial topology, every presheaf of sets on ${\mathscr{C}^{\mathrm{op}}}$ is a sheaf of sets. For a general Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$, the category ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$ of sheaves of sets on $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$ is a full subcategory of ${\mathscr{C}}\operatorname{-Set}$, the category of all presheaves of sets on ${\mathscr{C}^{\mathrm{op}}}$. The inclusion functor $\iota:{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})\to{\mathscr{C}}\operatorname{-Set}$ has a left adjoint $\sharp:{\mathscr{C}}\operatorname{-Set}\to{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$, called the sheafification functor. It is well known that the sheafification is a special type of localization; see [10, Proposition 1.3] or [15]. Therefore, ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$ is a localization of ${\mathscr{C}}\operatorname{-Set}$ by the multiplicative system of local isomorphisms [15] (or morphismes bicouvrants [2]).
Now we give definitions of ringed sites and sheaves of modules on ringed sites. For more details, please refer to [23, Chapter 18].
@@ 3.1. $J$-torsion modules
Denote by ${\mathscr{C}}_{x}$ the subcategory with the single object $x$ and the identity morphism $1_{x}$. The inclusion functor $\iota_{x}:{\mathscr{C}}_{x}\to{\mathscr{C}}$ induces a restriction functor $\textsl{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{x}\operatorname{-Mod}$ and its left adjoint $\textsl{ind}:\mathcal{O}_{x}\operatorname{-Mod}\to\mathcal{O}\operatorname{-Mod}$ is given by the left Kan extension. Explicitly, given an $\mathcal{O}_{x}$-module $W$ as well as an object $y$ in ${\mathscr{C}}$, one has
@@
Now let $P(x)=\textsl{ind}(\mathcal{O}_{x})$, and $\underline{S}$ be the submodule such that
@@
[…]
@@
[…]
@@
[…]
@@ 3.2. Torsion pairs induced by Grothendieck topologies
Trivial topology 00 $\mathcal{O}\operatorname{-Mod}$
@@
Maximal topology $\mathcal{O}\operatorname{-Mod}$ 00
@@ 4. A torsion theoretic characterization of sheaves
Throughout this subsection let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ and let $\mathcal{O}$ be a structure sheaf on the site $({\mathscr{C}^{\mathrm{op}}},J)$. Recall that an $\mathcal{O}$-module $V$ is right perpendicular to all $J$-torsion $\mathcal{O}$-modules in the sense of Geigle and Lenzing [12] if $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,V)=0=\operatorname{Ext}_{\mathcal{O}\operatorname{-Mod}}^{1}(T,V)$ for each $J$-torsion module $T$.
@@
$(2)\Leftrightarrow(3)$. In [7, Lemma 3.6] we have proved the conclusion for the special case that $J$ is the atomic Grothendieck topology. We use the same strategy to establish the equivalence for all Grothendieck topologies.
Suppose that $V$ is $J$-torsion free. It has an injective hull $E$ in $\mathcal{O}\operatorname{-Mod}$ since $\mathcal{O}\operatorname{-Mod}$ is a Grothendieck category. We claim that $E$ is torsion free. Otherwise, we have $\mathcal{T}_{J}(V)\neq 0$. But $V\cap\mathcal{T}_{J}(V)=0$ since $V$ is $J$-torsion free. This is impossible because the inclusion $V\to E$ is essential by [23, Definition 47.2.1]. Thus $E$ is $J$-torsion free and injective. Applying $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,-)$ to the short exact sequence $0\to V\to E\to E/V\to 0$ of $\mathcal{O}$-modules we deduce that $V$ is right perpendicular to every $J$-torsion module $T$ if and only if $E/V$ is $J$-torsion free. But this is equivalent to the condition that $V$ is $J$-saturated via applying the functor $\mathcal{T}_{J}$ to the short exact sequence.
$(1)\Rightarrow(2)$. Let $V$ be a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$. It is $J$-torsion free. Since the category of $J$-torsion modules is a localizing subcategory by [12, Proposition 2.5], applying [12, Proposition 2.2] we obtain a short exact sequence
@@
$(2)\Rightarrow(1)$: Suppose that $V$ is $J$-saturated. It is a separated presheaf of modules by Lemma 4.1. Take an arbitrary $x\in\operatorname{Ob}({\mathscr{C}})$, a certain $S\in J(x)$, and a matching family for $S$ of elements in $V$
@@
We need to show the existence of amalgamations. By [18, III.5, Lemma 2], the canonical map $\eta_{V}:V\to V^{+}$ is injective where $V^{+}=V^{\sharp}$ is the sheafification of $V$. Thus we obtain the following short exact sequence in $\mathcal{O}\operatorname{-Mod}$:
@@
Furthermore, by [18, III.5, Lemma 5], $V^{+}$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, and hence a $J$-saturated $\mathcal{O}$-module by the conclusion of the only if direction. Applying $\mathcal{T}_{J}$ to this short exact sequence, we deduce that $W$ is $J$-torsion free since $V$ is $J$-saturated.
@@
This result has been proved in [7, Corollary 3.9] for the atomic Grothendieck topology. Its proof still works in full generality. ∎
@@
The identification follows from Theorem 4.2  [^3] We shall remind the reader that the abelian structure on $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$ is not inherited from the abelian structure on $\mathcal{O}\operatorname{-Mod}$. Indeed, Given a monomorphism $\alpha:V\to W$ between two $J$-saturated modules, its cokernel in $\mathcal{O}\operatorname{-Mod}$ might not coincide with the cokernel in $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$., while the equivalence follows from [12, Proposition 2.2(d)]. ∎
Theorem 4.2 also provides a new description of the sheafification functor $\sharp$, which is more elementary compared to the ones using local isomorphisms or dense monomorphisms. This description has been observed for the atomic Grothendieck topology in [7, Section 3], and here we extend it to all Grothendieck topologies. Explicitly, given an $\mathcal{O}$-module $V$, the first step is to take the $J$-torsion free part $V_{F}=V/\mathcal{T}_{J}(V)$ and embeds it into an injective hull $I$ (which is $J$-saturated) to get a short exact sequence $0\to V_{F}\to I\to C\to 0$ of $\mathcal{O}$-modules. In the second step, we construct the pull back of $I\to C$ and $C_{T}=\mathcal{T}_{J}(C)\to C$ to obtain the following commutative diagram such that all rows and columns are exact:
@@
Then $V^{\sharp}\cong\widetilde{V}$. For more details, please refer to [7, Section 3].

### v3

@@ Abstract.
We prove that every Grothendieck topology induces a hereditary torsion pair in the category of presheaves of modules on a ringed site, and obtain a homological characterization of sheaves of modules: a presheaf of modules is a sheaf of modules if and only if it is saturated with respect to torsion presheaves, or equivalently, it is right perpendicular to torsion presheaves in the sense of Geigle and Lenzing. We also study Grothendieck topologies on directed categories ${\mathscr{C}}$ satisfying a certain locally finite condition, and show that every Grothendieck topology on ${\mathscr{C}}$ is a subcategory topology if and only if ${\mathscr{C}}$ is an artinian EI category. Consequently, in this case every sheaf category is equivalent to the presheaf category over a full subcategory of ${\mathscr{C}}$. Finally, we classify all Grothendieck topologies on a special type of noetherian EI categories, and extend many fundamental representation theoretic properties of ${\mathrm{FI}}$ and ${\mathrm{VI}}$ to their infinite full subcategories.
@@
Grothendieck topologies, rigid topologies, sheaves, torsion pairs, sheafification, sheaf cohomology, EI categories
Z.X. Di was partly supported by NSF of China (Grant No. 12471034), Scientific Research Fund of Fujian Province (Grant No. 605-52525002) and Scientific Research Fund of Huaqiao University (Grant No. 605-50Y22050); L.P. Li was partly supported by NSF of China (Grant No. 12171146); L. Liang was partly supported by NSF of China (Grant No. 12271230).
@@ 1. Introduction
Sheaves of modules on ringed sites have been widely applied in algebraic geometry, algebraic topology, and geometric representation theory. Recently, they were also used to investigate modular representation theory of finite groups (see for instances [3, 31, 32, 33]) and continuous representations of topological groups [8]. This new approach combines ideas and methods from topos theory, representation theory of categories, and representation theory of groups, and hence not only proposes a comprehensive technique to study representations of groups, but also reveals deep relations among distinct areas. A classical example is Artin’s theorem [23, III. 9, Theorems 1 and 2], from which it follows that the category of continuous representations of a topological group $G$ is equivalent to the category of sheaves of modules over a ringed site whose underlying category is an orbit category of $G$ equipped with the atomic Grothendieck topology.
@@
The following result gives a characterization of sheaves of modules over $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ in terms of the torsion theory mentioned above, and extends [8, Theorem 3.8] to full generality. Compared to the original definition of sheaves in terms of amalgamations of matching families, this characterization is homological, and is more elementary for applications in representation theory.
@@
It is direct to check that $J$-saturated $\mathcal{O}$-modules are precisely $\mathcal{O}$-modules right perpendicular to $J$-torsion $\mathcal{O}$-modules in the sense of Geigle-Lenzing [15]; for details, see [15, Section 1]. Consequently, applying [15, Propositions 2.2 and 2.5], one immediately deduces the following result:
@@
It has been pointed out in [2, IV.9] by Grothendieck and Verdier that every Grothendick topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid (or called a subcategory topology) when ${\mathscr{C}}$ is a finite Karoubi category, and in this case ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is equivalent to the category of presheaves of modules on a full subcategory of ${\mathscr{C}^{\mathrm{op}}}$. Similar results have been established by Hemelaer [18] and Lindenhovius [22] for posets satisfying certain extra conditions. Although it seems to be a very difficult task to classify small categories whose Grothendieck topologies are rigid, in this paper we obtain a satisfactory answer for categories satisfying the following combinatorial condition: the binary relation $\leqslant$ on $\operatorname{Ob}({\mathscr{C}})$ defined by setting $x\leqslant y$ if and only if ${\mathscr{C}}(x,y)\neq\emptyset$ is a partial order. We call them directed categories. The reason we choose them is plain: many important categories of particular interest in representation theory are directed categories, for examples, posets, orbit categories of groups, quivers without loops or oriented cycles, and etc.
@@
Let ${\mathscr{C}}$ be a directed category such that ${\mathscr{C}}(x,x)$ is finite for every $x\in\operatorname{Ob}({\mathscr{C}})$. Then every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if ${\mathscr{C}}$ is a noetherian EI category; that is, the poset $(\operatorname{Ob}({\mathscr{C}}),\leqslant)$ satisfies the ascending chain condition and every endomorphism in ${\mathscr{C}}$ is an isomorphism.
Remark 1.5.
The if direction actually holds without the assumption that every ${\mathscr{C}}(x,x)$ is a finite set; see Remarks 5.10 and 5.14.
Consequently, when ${\mathscr{C}}$ is a noetherian EI category, questions concerning ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ can be completely transferred to questions about representations of full subcategories of ${\mathscr{C}}$, so we have no “generic” sheaf theory in this situation. Explicitly, by the Comparison Lemma [17, C2.2, Theorem 2.2.3] (see also [17, C2.2, Example 2.2.4(d)]), one has:
@@
However, when ${\mathscr{C}}$ is not noetherian, a classification of Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ becomes much harder. In this paper we solve this question only for EI categories ${\mathscr{C}}$ of type $\mathbb{N}$ (see the beginning of Section 6 for a precise definition) or type $\mathbb{Z}$. Although it seems to the reader that too many combinatorial restrictions are imposed on them, there are still quite a few interesting examples whose representation theory are actively studied (see for instances [6, 7, 25, 26, 27, 28]):
  • skeletal full subcategories of the category ${\mathrm{FI}}$ of finite sets and injections, where ${\mathrm{FI}}$ is equivalent to the opposite category of an orbit category of the permutation group on ${\mathbb{N}}$ ([17, Examples A2.1.11(h) and D3.4.10]);
  • skeletal full subcategories of the opposite category of the category $\mathrm{FS}$ of finite sets and surjections, where $\mathrm{FS}$ is equivalent to an orbit category of the automorphism group of the free Boolean algebra on a countable infinity of generators, or the self-homeomorphism group of the Cantor set ([17, Example D3.4.12]);
@@
We can use sheaf theory to deduce some interesting results in representation theory of combinatorial categories. For instance, let $k$ be a field of characteristic 0, and let ${\mathscr{C}}$ be a skeleton subcategory of the category ${\mathrm{FI}}$ of finite sets and injections, or the category ${\mathrm{VI}}_{q}$ of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved by Gan-Li [12] (see also Sam-Snowden [27] for ${\mathrm{FI}}$) that every indecomposable projective ${\mathscr{C}}$-module over $k$ is also injective. Surprisingly, this locally self-injective property and many other fundamental results of ${\mathscr{C}}$-modules established by Church-Ellenberg-Farb-Nagpal [7], Gan-Li [12], Gan-Li-Xi [14], Nagpal [25, 26], and Sam-Snowden [27, 28] extend to representations of its infinite full subcategories.
@@
Let ${\mathscr{C}}$ be specified as above, $k$ a commutative noetherian ring, and ${\mathscr{D}}$ a full subcategory of ${\mathscr{C}}$ containing infinitely many objects. Then every finitely generated ${\mathscr{D}}$-module over $k$ is noetherian. Moreover, if $k$ is a field of characteristic 0, then one has
  1. (1)
    every indecomposable projective ${\mathscr{D}}$-module over $k$ is also injective;
  2. (2)
    the category ${\mathscr{D}}\operatorname{-mod}$ of finitely generated ${\mathscr{D}}$-modules has enough injectives;
  3. (3)
    up to isomorphism, an indecomposable injective ${\mathscr{D}}$-module is one of the following three:
    • a summand of $k{\mathscr{D}}(x,-)$ for a certain $x\in\operatorname{Ob}({\mathscr{D}})$,
    • a summand of the restriction of $k{\mathscr{C}}(x,-)$ for a certain $x\in\operatorname{Ob}({\mathscr{C}})\setminus\operatorname{Ob}({\mathscr{D}})$,
    • a finite dimensional injective ${\mathscr{D}}$-module;
  4. (4)
    every finitely generated ${\mathscr{D}}$-module has finite injective dimension;
  5. (5)
    one has the following equivalences
    ${\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-fdmod}\simeq{\mathscr{C}}\operatorname{-fdmod},$
    where ${\mathscr{D}}\operatorname{-fdmod}$ (resp., ${\mathscr{C}}\operatorname{-fdmod}$) is the category of finite dimensional ${\mathscr{D}}$-modules (resp, ${\mathscr{C}}$-modules), and ${\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-fdmod}$ is the Serre quotient.
Remark 1.9.
A standard tool to prove these results for ${\mathscr{C}}$ is the shift functor introduced by Church-Ellenberg-Farb [6] for representations of ${\mathrm{FI}}$ and Nagpal [25] for representations of ${\mathrm{VI}}_{q}$. Unfortunately, it does not work in this much wider situation as the objects in ${\mathscr{D}}$ in general do not occur in a periodic way, which is essential for people to define the shift functor. However, using sheaf theory, one can deduce this theorem via an almost transparent way from those established results of ${\mathscr{C}}$.
@@ 2.1. Grothendieck topologies
Throughout this paper we fix $J$ to be a rule assigning to each object $x$ a collection $J(x)$ of left ideals of morphisms on $x$. The following definition is taken from [23, III.2, Definition 1], which defines a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ if $J$ satisfies all three axioms. Here we use the dual version, since we mainly consider representations of ${\mathscr{C}}$ which are covariant functors, rather than presheaves which are contravariant functors.
@@
The following fact is well known; see [23, Section III.2]. For the convenience of the reader, we give a complete proof.
@@ 2.2. Sheaf theory
Fix a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ and let $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ be the corresponded site. The following definitions are taken from [23, III.4].
@@
When $J$ is the trivial topology, every presheaf of sets on ${\mathscr{C}^{\mathrm{op}}}$ is a sheaf of sets. For a general Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$, the category ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$ of sheaves of sets on $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$ is a full subcategory of ${\mathscr{C}}\operatorname{-Set}$, the category of all presheaves of sets on ${\mathscr{C}^{\mathrm{op}}}$. The inclusion functor $\iota:{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})\to{\mathscr{C}}\operatorname{-Set}$ has a left adjoint $\sharp:{\mathscr{C}}\operatorname{-Set}\to{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$, called the sheafification functor. It is well known that the sheafification is a special type of localization; see [11, Proposition 1.3] or [19]. Therefore, ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$ is a localization of ${\mathscr{C}}\operatorname{-Set}$ by the multiplicative system of local isomorphisms [19] (or morphismes bicouvrants [2]).
Now we give definitions of ringed sites and sheaves of modules on ringed sites. For more details, please refer to [29, Chapter 18].
@@ 3.1. $J$-torsion modules
Denote by ${\mathscr{C}}_{x}$ the subcategory with the single object $x$ and the identity morphism $1_{x}$. The inclusion functor $\iota_{x}:{\mathscr{C}}_{x}\to{\mathscr{C}}$ induces a restriction functor $\operatorname{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{x}\operatorname{-Mod}$ and its left adjoint $\operatorname{ind}:\mathcal{O}_{x}\operatorname{-Mod}\to\mathcal{O}\operatorname{-Mod}$ is given by the left Kan extension. Explicitly, given an $\mathcal{O}_{x}$-module $W$ as well as an object $y$ in ${\mathscr{C}}$, one has
@@
Now let $P(x)=\operatorname{ind}(\mathcal{O}_{x})$, and $\underline{S}$ be the submodule such that
@@
which is clearly a left ideal of morphisms on $x$. For a nonempty full subcategory $\mathfrak{A}$ of $\mathcal{O}\operatorname{-Mod}$, we define $\mathcal{A}(\mathfrak{A})$ to be the rule assigning to each $x\in\operatorname{Ob}({\mathscr{C}})$ the collection
@@
Clearly, $\mathcal{A}(\mathfrak{A})$ satisfies the maximal axiom. It also satisfies the stability axiom.
@@
The rule $\mathcal{A}(\mathfrak{A})$ satisfies the stability axiom.
@@ 3.2. Torsion pairs induced by Grothendieck topologies
Trivial topology 0 $\mathcal{O}\operatorname{-Mod}$
@@
Maximal topology $\mathcal{O}\operatorname{-Mod}$ 0
@@ 4. A torsion theoretic characterization of sheaves
Throughout this subsection let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ and let $\mathcal{O}$ be a structure sheaf on the site $({\mathscr{C}^{\mathrm{op}}},J)$. Recall that an $\mathcal{O}$-module $V$ is right perpendicular to all $J$-torsion $\mathcal{O}$-modules in the sense of Geigle-Lenzing [15] if $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,V)=0=\operatorname{Ext}_{\mathcal{O}\operatorname{-Mod}}^{1}(T,V)$ for each $J$-torsion module $T$.
@@
(2)(3)(2$\Leftrightarrow$(3)( 2 ) ⇔ ( 3 ). In [8, Lemma 3.6] we have proved the conclusion for the special case that $J$ is the atomic Grothendieck topology. We use the same strategy to establish the equivalence for all Grothendieck topologies.
Suppose that $V$ is $J$-torsion free. It has an injective hull $E$ in $\mathcal{O}\operatorname{-Mod}$ since $\mathcal{O}\operatorname{-Mod}$ is a Grothendieck category. We claim that $E$ is torsion free. Otherwise, we have $\mathcal{T}_{J}(V)\neq 0$. But $V\cap\mathcal{T}_{J}(V)=0$ since $V$ is $J$-torsion free. This is impossible because the inclusion $V\to E$ is essential by [29, Definition 47.2.1]. Thus $E$ is $J$-torsion free and injective. Applying $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,-)$ to the short exact sequence $0\to V\to E\to E/V\to 0$ of $\mathcal{O}$-modules we deduce that $V$ is right perpendicular to every $J$-torsion module $T$ if and only if $E/V$ is $J$-torsion free. But this is equivalent to the condition that $V$ is $J$-saturated via applying the functor $\mathcal{T}_{J}$ to the short exact sequence.
(1)(2)(1$\Rightarrow$(2)( 1 ) ⇒ ( 2 ). Let $V$ be a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$. It is $J$-torsion free. Since the category of $J$-torsion modules is a localizing subcategory by [15, Proposition 2.5], applying [15, Proposition 2.2] we obtain a short exact sequence
@@
(2)(1)(2$\Rightarrow$(1)( 2 ) ⇒ ( 1 ): Suppose that $V$ is $J$-saturated. It is a separated presheaf of modules by Lemma 4.1. Take an arbitrary $x\in\operatorname{Ob}({\mathscr{C}})$, a certain $S\in J(x)$, and a matching family for $S$ of elements in $V$
@@
We need to show the existence of amalgamations. By [23, III.5, Lemma 2], the canonical map $\eta_{V}:V\to V^{+}$ is injective where $V^{+}=V^{\sharp}$ is the sheafification of $V$. Thus we obtain the following short exact sequence in $\mathcal{O}\operatorname{-Mod}$:
@@
Furthermore, by [23, III.5, Lemma 5], $V^{+}$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, and hence a $J$-saturated $\mathcal{O}$-module by the conclusion of the only if direction. Applying $\mathcal{T}_{J}$ to this short exact sequence, we deduce that $W$ is $J$-torsion free since $V$ is $J$-saturated.
@@
This result has been proved in [8, Corollary 3.9] for the atomic Grothendieck topology. Its proof still works in full generality. ∎
@@
The identification follows from Theorem 4.2  [^3] We shall remind the reader that the abelian structure on $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$ is not inherited from the abelian structure on $\mathcal{O}\operatorname{-Mod}$. Indeed, Given a monomorphism $\alpha:V\to W$ between two $J$-saturated modules, its cokernel in $\mathcal{O}\operatorname{-Mod}$ might not coincide with the cokernel in $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$., while the equivalence follows from [15, Proposition 2.2(d)]. ∎
Theorem 4.2 also provides a new description of the sheafification functor $\sharp$, which is more elementary compared to the ones using local isomorphisms or dense monomorphisms. This description has been observed for the atomic Grothendieck topology in [8, Section 3], and here we extend it to all Grothendieck topologies. Explicitly, given an $\mathcal{O}$-module $V$, the first step is to take the $J$-torsion free part $V_{F}=V/\mathcal{T}_{J}(V)$ and embeds it into an injective hull $I$ (which is $J$-saturated) to get a short exact sequence $0\to V_{F}\to I\to C\to 0$ of $\mathcal{O}$-modules. In the second step, we construct the pull back of $I\to C$ and $C_{T}=\mathcal{T}_{J}(C)\to C$ to obtain the following commutative diagram such that all rows and columns are exact:
@@
Then $V^{\sharp}\cong\widetilde{V}$. For more details, please refer to [8, Section 3].